OpenStudy (anonymous):

find the average rate of change of F(x)=sec(x) from 0 to pi/4

OpenStudy (aum):

Average rate of change of \(f(x) = \sec(x)\) from x = 0 to x = \(\pi/4\) is: \(\Large \frac {f(\pi/4)-f(0)}{\pi/4 - 0} = \frac {\sec(\pi/4)-\sec(0)}{\pi/4} = ?\)